Suppose an electron has a wavelength of 1.25 m.

What is the velocity of the electron in meters per second?
Answer #1

Solution)

Given,

Wavelength, lambda=1.25 m

We know, velicity, v= h/m*lambda

V=(6.626*10^-34)/(9.1*10^-31)*(1.25)=0.000582 m/s (Ans)

============